OpenStudy (anonymous):

What is the value of 54 – 4a2 + 3b3 when a = −2 and b = 4? for this i keep getting 106 im correct or wrong

OpenStudy (solomonzelman):

incorrect, sorry

OpenStudy (anonymous):

\[54-4(-2)^2+3(4)^3\] and a calculator

OpenStudy (solomonzelman):

and that is not 106, I know from wolfram

OpenStudy (anonymous):

i got 262

OpenStudy (anonymous):

i didn't

OpenStudy (solomonzelman):

nope, don't think so

OpenStudy (anonymous):

i meant 230

OpenStudy (anonymous):

so what im I doing wrong

OpenStudy (texaschic101):

54 - 4(-2^2) + 3(4^3) = 54 - 4(4) + 3(64) = try it now

OpenStudy (anonymous):

54-4*(-2)*24*(-2)=-8. -8*2=-16+3*4*3

OpenStudy (anonymous):

yea so it is 230

OpenStudy (anonymous):

but i think ur math is wrong lorenav

OpenStudy (anonymous):

u eventually git 54 -16 +192

OpenStudy (texaschic101):

yep...230 54 - 4(4) + 3(64) = 54 - 16 + 192 246 - 16 = 230

OpenStudy (anonymous):

yes I saw thank you guys I appreciate your help im not smart at all but i always try.

OpenStudy (anonymous):

lol np

OpenStudy (anonymous):

thank you Da_Homie and Texaschic101 for the explanation

OpenStudy (texaschic101):

no problem...happy to help :)